OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)?

OpenStudy (anonymous):

y = 2x – 8 y = – 2x + 9 y = 2x – 11 y = – 2x + 6

OpenStudy (anonymous):

Could it be B?

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(2\quad ,&-7)\quad &(5\quad ,&-1) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)

jigglypuff314 (jigglypuff314):

it would not be B first you would want to find the slope given the formula that jdoe0001 gave you plug x1 = 2 , y1 =-7 , x2 = 5 , y2 = -1 so \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } = \frac{ (-1)-(-7) }{ (5)-(2) }\]

jigglypuff314 (jigglypuff314):

first find the slope, tell me what you got :)

OpenStudy (anonymous):

6/3 or 2/1

jigglypuff314 (jigglypuff314):

yep and the equations are in this format y = mx + b when m = slope and b = y-intercept so since slope = 2 = m y = 2x + b ^ then the answer can not be B or D

OpenStudy (anonymous):

So A?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

are you just guessing? -.-" no, it's not A...

OpenStudy (anonymous):

Frankly, yes. I've been doing this since early morning and I'm just trying to get it done. Thanks for your help though